OpenStudy (anonymous):

Use trig substitution to solve the integral of 1/(4x^2(1-x^2)^(1/2)

hartnn (hartnn):

so what substitution are you thinking ? any thoughts ?

OpenStudy (anonymous):

I'm pretty sure you use \[x=asin (\Theta)\]

OpenStudy (anonymous):

and a=1

hartnn (hartnn):

correct. tried it ?

OpenStudy (anonymous):

so x=sin(theta) and dx=cos(theta) d(theta)

hartnn (hartnn):

yup, what about 1-x^2 = .. ?

hartnn (hartnn):

and (1-x^2)^1/2 =... ?

OpenStudy (anonymous):

Looking at my work I got that that is cos(theta)

hartnn (hartnn):

correct, so what does your integral turn into now ? everything in terms of theta

OpenStudy (anonymous):

So then I plugged everything in and got \[\frac{ 1 }{ 4 }\int\limits_{}^{}\frac{ 1 }{ \sin ^{2}(\Theta)-\cos(\Theta) } d(\Theta)\]

hartnn (hartnn):

4x^2 (1-x^2) would be 4 (sin theta)^2 cos theta right ? where did u get the - *minus* sign there ?

hartnn (hartnn):

also your numerator is incorrect..

OpenStudy (anonymous):

I don't know why I typed that minus sign.

OpenStudy (anonymous):

I didn't get it. I just have sin^2cos

hartnn (hartnn):

it just shouldn't be there. ... about the numerator, the dx becomes dx = cos theta d theta you missed cos theta term...

OpenStudy (anonymous):

I sure did

OpenStudy (anonymous):

so there should be a cos(theta) as the numerator

hartnn (hartnn):

yup, which gets cancelled with the denominator cos theta so you are left with just integral 1/ sin^2 theta dtheta ok til here ?

OpenStudy (anonymous):

Yep, got it

hartnn (hartnn):

can u continue ?

OpenStudy (anonymous):

couldn't you do (1/sin^2(theta) = csc^2(theta)?

hartnn (hartnn):

absolutely, you can :)

OpenStudy (anonymous):

then use the identity csc^2(theta) = -cot(theta)

hartnn (hartnn):

don't forget the +c and then to resubstitute theta = sin^-1 x

OpenStudy (anonymous):

theta = sin^-1x?

hartnn (hartnn):

yes! x = sin theta so, theta = sin^-1 x inverse sin function

OpenStudy (anonymous):

Ah, okay

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }(-\cot(\sin^{-1} (x))?\]

hartnn (hartnn):

yeah, i get the same, just with +c in the end

OpenStudy (anonymous):

haha, sorry :P

OpenStudy (anonymous):

I remember you have to draw a triangle at one point and solve for the missing side?

hartnn (hartnn):

oh, if you need an algebraic answer instead of trigo one so, draw a right triangle first...

OpenStudy (anonymous):

oh, okay. I think I remember now. Then sin(theta)=x/1

OpenStudy (anonymous):

so then you just plug in the values and solve

hartnn (hartnn):

right! correct way...if you get stuck, ask me

OpenStudy (anonymous):

okay, awesome. Thank you!

hartnn (hartnn):

welcome ^_^